# AARK: An Open Toolkit for Autonomous Racing Research

### v1

@@ Abstract
Autonomous racing demands safe control of vehicles at their physical limits for extended periods of time, providing insights into advanced vehicle safety systems which increasingly rely on intervention provided by vehicle autonomy. Participation in this field carries with it a high barrier to entry. Physical platforms and their associated sensor suites require large capital outlays before any demonstrable progress can be made. Simulators allow researches to develop soft autonomous systems without purchasing a platform. However, currently available simulators lack visual and dynamic fidelity, can still be expensive to buy, lack customisation, and are difficult to use. AARK provides three packages, ACI, ACDG, and ACMPC. These packages enable research into autonomous control systems in the demanding environment of racing to bring more people into the field and improve reproducibility: ACI provides researchers with a computer vision-friendly interface to Assetto Corsa for convenient comparison and evaluation of autonomous control solutions; ACDG enables generation of depth, normal and semantic segmentation data for training computer vision models to use in perception systems; and ACMPC gives newcomers to the field a modular full-stack autonomous control solution, capable of controlling vehicles to build from. AARK aims to unify and democratise research into a field critical to providing safer roads and trusted autonomous systems.
@@ I INTRODUCTION
When vehicles are operated at the edges of their physical limitations a balance must be struck between operator and platform to ensure a favourable outcome; the slightest misstep on either part leads to catastrophic outcomes for both. This place, where a platform is pushed to its absolute capability, is where vehicle operators find themselves before incidents that result in damage, injury, and loss of life. It is therefore imperative to explore vehicle behaviour in these situations to improve road safety. The use of automated assistive systems to bring vehicles back from the brink, in-spite of its operator, are becoming increasingly common. Traction control, electronic stability control, and anti-lock braking are all automatic systems that respond to and remedy a vehicle’s rapid approach toward or surpassing of its capabilities. These systems rely on making decisions in response to intrinsic vehicle sensors that influence vehicle behaviour in a limited form of autonomy that usurps control from the operator. Advanced Driver Assistance Systems (ADAS) takes this further. ADAS exists in partnership with the operator, using extrinsic sensors to respond to the environment external to the vehicle. This manifests as adaptive cruise control, lane following, blind spot alerts and automatic braking to increase safety. In this way, manufacturers have been expanding the level of autonomy present in vehicles to become trusted partners of drivers. So far this expansion of autonomy in production vehicles has built up from intrinsic to extrinsic assistive autonomy. Conversely, autonomous racing provides a top-down approach, where fully autonomous systems, capable of controlling vehicles on their own, can be used to provide assurances of trust and efficacy for autonomous systems in general, while providing key insights into how autonomy in current offerings can be expanded to provide safer roads for everyone. Traditionally, motorsport has been used to provide an environment that forces vehicles to be operated at the limits of their capability, revealing their durability, stability and performance. In such events, for a driver to be competitive they must continually flirt with and even cross the barrier between being in and out of control. A similar statement can be made when replacing drivers, in such events, with autonomous systems. This expands the concerns of autonomous driving of road vehicles by providing an imperative that ensures agents push vehicles to their limits. Therefore, systems that race successfully will have an implicit comfort with situations human operators find themselves in prior to disaster; furthering understanding into autonomy enabled safety systems.
@@ II RELATED WORK
Previous efforts to provide a simulator for autonomous racing have failed to proliferate into consistent use across research institutions; preventing the unification of the field and limiting reproducibility. Factors preventing this are specific to each simulator. In general, they have been complicated to use, short on customisation options, closed source, or lack visual and dynamic fidelity. A series of pedagogical platforms have been developed to support educational programs such as AWS DeepRacer [1] and F1TENTH [2, 3]. Despite being easy to use, these platforms lack modelling rigour and limit customisation of vehicles. A similar deviation from reality in their approach to modelling is present in Open AI Gym’s Car Racing [4] and Track Mania [5]. Collectively, these simulations use RC scale or imagined vehicles without veracious dynamics modelling. Similar deficits in simulator fidelity are also present in full-scale vehicle simulators. Car Learning to Act (CARLA) [6], Generalized RAcing Intelligence Competition (GRAIC) [7, 8], Autoware Simulator (AWSIM) and Learn-to-Race (L2R) are all simulators that have been used in full-scale autonomous racing challenges. All of these examples are modified versions of simulators built primarily for the evaluation of driverless road vehicles. Although autonomous racing and driving appear to be analogues of one another, there are a series of simplifications that can be applied to dynamics simulation when vehicles are intended to be driven well within their physical limits. Self-driving vehicle simulators neglect high fidelity simulation of aerodynamics, fuel usage, tyre degradation and suspension geometry. This is because these have a negligible effect for road vehicles driven under normal conditions. Additionally, some of these simulators provide visually sparse environments that limit the application of computer vision techniques as part of an autonomous control stack.
@@
In contrast, Assetto Corsa (AC) is a customisable, racing-specific simulator that has a large collection of virtualised world-renowned race tracks and cars rendered to a high level of visual quality. AC’s inherent customisablity allows it to be used to simulate existing real-world autonomous racing leagues: Indy Autonomous Challenge [9], Autonomous Karting Series [10], Roborace [11] and A2RL [12]. Although the simulator itself is closed source, a mature modification community exists which have exposed internal methods of customisation and developed a range of additional assets such as vehicles, circuits, and visual improvements. AC is also regularly on sale for \$4USD [13] making it significantly more accessible compared to proprietary simulators used in motorsport. The desirability of AC has also been validated by its use in other work. In  [14] there is a section dedicated to validation of their control solution using AC and in the future, an interface would be published openly; this has yet to happen. Most recently a pre-print [15] developed concurrently to AARK was uploaded to arxiv detailing a similar interface for AC. As this interfacefocuses on control, the interfaces can not provide agents with visual sensor information or simulated INS readings, cannot be used to generate machine vision data, and they do not provide a modular full stack autonomous control solution.
@@
Enter the Adelaide Autonomous Racing Kit (AARK); an open source toolkit that enables research, development and validation of autonomous systems capable of operating vehicles pushed to their limits. AARK represents a feature rich, extensible manifestation of autonomous racing researcher’s desire to validate and compare autonomous systems in AC. Built around AC, AARK offers researchers a platform capable of generating data to train machine learning systems in addition to validating and comparing autonomous systems via a series of packages:
@@
  • AC Interface (ACI) - A harness that enables recording human drivers and enables autonomous systems to receive observations from and control vehicles in AC
@@
The source code to each of these packages is available on Github https://github.com/Adelaide-Autonomous-Racing-Challenge with accompanying documentation https://www.adelaideautonomous.racing/.
@@ III ADELAIDE AUTONOMOUS RACING KIT
The Adelaide Autonomous Racing Kit (AARK) positions itself in the no man’s land between two separate philosophical approaches to the autonomous racing problem; classical control and reinforcement learning (RL). Classical approaches to autonomous racing systems are built up from modular components that each handle specific tasks. This leads to systems that require laborious hand tuning of non-linear optimisation parameters which are specific to each particular vehicle and circuit [16, 17, 14, 18]. Conversely, RL approaches ingest environmental state and directly produce control outputs. In the current literature, classical autonomous systems heavily rely on known pose, GPS, LiDAR, and vehicle dynamics modelling. By contrast, previous RL methods have oracle knowledge of the simulation’s internal state [19, 20, 21] and have the luxury of acting synchronously with their environment [22, 1]. For RL approaches to become practically useful, they need to evolve from relying on perfect state information and account for the fundamentally asynchronous nature of reality. The strengths of both approaches have the potential to address the weaknesses of the other. In fact, machine learning approaches are being integrated into classic autonomous system’s modules [23], albeit slowly over time. For these fields to unify and develop together they require a common environment that is equally convenient for researchers of both methodologies to use. AARK has the potential to bridge this gap by providing a development and evaluation platform for autonomous racing systems that cater to both of these disciplines.
@@ IV ASSETTO CORSA INTERFACE
At its core AARK provides an interface between the popular racing simulator AC and autonomous agents, Figure 1. The enabling package, ACI, achieves this by scraping vehicle state from the simulator’s memory, capturing frames displayed using FFMPEG and relaying control inputs to AC via a virtual controller. At a frequency of 1̃00Hz, ACI outputs 190 different readings describing the simulated vehicle’s current state. These readings contain detailed information regarding the vehicle’s ground truth state, intended for evaluation and debugging. Accompanying this are 17 readings from a simulated Inertial Navigation System (INS), intended to be passed to the autonomous driver. The INS provides accelerometer, gyroscope, odometer and GPS readings with artificial noise and drift to better reflect real sensors. Although it is not our intention, the ground truth state can be passed to the autonomous driver to baseline against previous RL techniques. With these readings, ACI also provides image frames captured of the simulator’s application window. Frame captures are non-blocking, allowing systems to continue performing operations that depend on sensor readings while waiting for new visual input. The frequency of frame captures is 30Hz by default but can be configured by the user to suite their requirements and computer’s performance. Once an observation has been processed by the autonomous system it returns $\text{throttle},\text{brake},\text{steering}][ throttle , brake , steering$ ] to ACI which represents the inputs that should be applied to the vehicle.
Additionally, ACI provides two methods to capture driver data for analysis or training. Leveraging an AC plugin, AC Telemetry Interface (ACTI), vehicle performance data can be captured and written out in the prolific MoTeC format. MoTeC data is currently used to analyse both vehicle and driver performance across motorsport, providing access to a suite of ready-made visualisation tools to verify system behaviour. Vast libraries of driver MoTeC data exist online, for a wide range of operator proficiencies, to compare autonomous systems to human drivers. However, MoTeC does not offer a convenient way to package visual information synchronised with state information in the necessary form for training AI-based approaches. ACI addresses this deficit by recording sessions as a series of image, state pairs at a frequency of  30Hz. This format of recording can be used to train machine vision systems with either directly or indirectly via post-processed data produced by AC Data Generator (ACDG). More details on the implementation framework for ACI are given in appendix VIII-A.
@@ V DATA GENERATOR
Machine learning has demonstrated efficacy in tasks that estimate the properties of objects from images of those objects. Many of these properties can be leveraged to provide information to an autonomous system about its environment, such as depth, surface normal vectors and semantics. These sources of information have been under utilised in previous approaches that rely heavily on LiDAR, raw camera feeds, or ego state. The barrier, and likely reason for their lack of proliferation, to applying machine learning based vision techniques is their requirement for a large body of domain specific annotated data. For example, to train a semantic segmentation model for use in autonomous racing, video data from each different race track needs to be gathered and human annotators need to specify which pixels in each image belong to which racing specific semantic class. ACDG automatically generates multiple types of annotated machine learning data from recordings gathered by ACI. Figure 2 shows an example of semantic, depth, and surface normal annotations generated for an image by ACDG. This is achieved by; recreating the recorded camera perspective at each time step, casting rays, and extracting information about where those rays intersect with the circuit’s mesh. Depth and surface normal annotations can be generated for any circuit.
ACDG’s initial release contains semantic mappings for eight circuits, shown in Figure 4. These include six official AC maps: Imola, Circuit de Spa-Francorchamps, Autodromo Nazionale di Monza, Silverstone, Vallenlunga, Nordschleife, in addition to two community-made tracks: Mount Panorama and Yas Marina. Tracks were chosen based on their respective challenge to drivers and agents, alike, with varying levels complexity. The complexity of a track ranges from Monza, which has many long straights and very few corners, to Nurburgring - one of the most challenging tracks in the world, with 170 turns spaced over 25km. Mt Panorama and Spa features large changes in elevation and blind corners. Silverstone and Monza contain several corner elements that need to be connected to prioritise exit speed on to straights rather than speed through corners. Vallelunga focuses on extremely technical slow speed corner with hairpin style loops. Yas Marina and Imola are both varied and complex circuits in their own right which are included to support parity with the real-world autonomous racing league A2RL [12].
Each of these race tracks includes a semantic segmentation dataset and a pre-trained deep learning model for use in researchers’ control solutions. If researchers wish to develop a dataset for a not yet supported circuit, a series of video tutorials demonstrating how to add semantic segmentation data generation support for new tracks are available here: https://www.adelaideautonomous.racing//docs/acdg/getting-started/tutorials/.
@@ V-A DATASETS
To accelerate researcher’s ability to participate in autonomous racing, we release a series of datasets created using AARK which are available here: https://www.adelaideautonomous.racing/downloads/.
@@ V-A1 Behaviour
For each of the supported racetracks associated recordings of amateur human control were gathered. Roughly eight laps of human control is available for use in both the ACI and MoTeC formats. MoTeC is compatible with i2, a free data analysis software, used by professional motorsport teams to analyse driver and vehicle performance. By collecting telemetry data in this format, autonomously controlled vehicle behaviour can be compared against their human driven counter parts. This can provide insights into how autonomous control differs from race car drivers and vice versa. Recordings provided in the ACI format are compatible with behavioural cloning or supervised machine learning methods. Additionally, the ACI format can be used to generate a series of machine vision data for deep learnt perception systems.

### v2

@@ Abstract
Autonomous racing demands safe control of vehicles at their physical limits for extended periods of time, providing insights into advanced vehicle safety systems which increasingly rely on intervention provided by vehicle autonomy. Participation in this field carries with it a high barrier to entry. Physical platforms and their associated sensor suites require large capital outlays before any demonstrable progress can be made. Simulators allow researchers to develop soft autonomous systems without purchasing a platform. However, currently available simulators lack visual and dynamic fidelity, can be expensive to license, lack customisation, and are difficult to use. AARK provides four packages, ACI, ACDG, ACMPC and ACRL. These packages enable research into autonomous control systems in the demanding environment of racing to bring more people into the field and improve reproducibility: ACI provides researchers with a computer vision-friendly interface to Assetto Corsa for convenient comparison and evaluation of autonomous control solutions; ACDG enables generation of depth, normal and semantic segmentation data for training computer vision models to use in perception systems; ACMPC and ACRL provide full stack autonomous, and reinforcement learning based solutions capable of controlling vehicles to build from. AARK aims to unify and democratise research into a field critical to providing safer roads and trusted autonomous systems.
@@ I INTRODUCTION
When vehicles are operated at the edges of their physical limitations a balance must be struck between operator and platform to ensure a favourable outcome; the slightest misstep on either part leads to catastrophic outcomes for both. This place, where a platform is pushed to its absolute capability, is where vehicle operators find themselves before incidents that result in damage, injury, and loss of life. It is therefore imperative to explore vehicle behaviour in these situations to improve road safety. The use of automated assistive systems to bring vehicles back from the brink, in-spite of its operator, are becoming increasingly common. Traction control, electronic stability control, and anti-lock braking are all automatic systems that respond to and remedy a vehicle’s rapid approach toward or surpassing its capabilities. These systems rely on making decisions in response to intrinsic vehicle sensors, in a limited form of autonomy that usurps control from the operator. Advanced Driver Assistance Systems (ADAS) takes this further. ADAS exists in partnership with the operator, using extrinsic sensors to respond to the environment external to the vehicle. This manifests as adaptive cruise control, lane following, blind spot monitoring and automatic braking. In this way, manufacturers have been expanding the level of autonomy present in vehicles to become trusted partners of drivers. At the other end of the spectrum, fully autonomous systems, capable of controlling vehicles on their own, can be used to provide insights into how current assistive offerings can be expanded further to provide safer roads. However, in the classical self-driving problem situations where the vehicle is exerting safety critical control at the edge of platform capability are rare. These events are exactly what occur before many road incidents. When their physical limits are pushed vehicles behave drastically different from when they are use under normal operation. Inputs become detached to their outputs, causing road users to make inputs that push the vehicle further out of control, due to misaligned expectations. Without validation of autonomy’s capacity to deal with this paradigm shift in control dynamics, how can they be trusted to safely negotiate and recover vehicles from the brink of disaster? Traditionally, motorsport has been used to provide an environment that forces vehicles to be operated at the limits of their capability, revealing durability, stability and performance. In such events, for a driver to be competitive they must continually flirt with and even cross the barrier between being in and out of control. A similar statement can be made when replacing drivers, in such events, with autonomous systems. This expands the concerns of traditional self-driving autonomy by providing an imperative that ensures agents operate vehicles at the limits of physical capability. Control solutions that are able to race effectively and safely will have an implicit ability to exert corrective control in situations where human operators find themselves prior to disaster.
@@ II RELATED WORK
Previous efforts to provide a simulator for autonomous racing have failed to proliferate into consistent use across research institutions; preventing the unification of the field and limiting reproducibility. Preventing factors are specific to each simulator. In general, they have been complicated to use, short on customisation options, closed source, or lack visual and dynamic fidelity. A series of pedagogical platforms have been developed to support educational programs such as AWS DeepRacer [1] and F1TENTH [2, 3]. Despite being easy to use, these platforms lack modelling rigour and limit customisation of vehicles. A similar deviation from reality in their approach to modelling is present in Open AI Gym’s Car Racing [4] and Track Mania [5]. Collectively, these simulations use RC scale or imagined vehicles without veracious dynamics modelling. Similar deficits in simulator fidelity are also present in full-scale vehicle simulators. Car Learning to Act (CARLA) [6], Generalized RAcing Intelligence Competition (GRAIC) [7, 8], Autoware Simulator (AWSIM) and Learn-to-Race (L2R) are all simulators that have been used in full-scale autonomous racing challenges. All of these examples are modified versions of simulators built primarily for the evaluation of driverless road vehicles. Although autonomous racing and driving appear to be analogues of one another, there are a series of simplifications that can be applied to dynamics simulation when vehicles are intended to be driven well within their physical limits. Self-driving vehicle simulators neglect high fidelity simulation of aerodynamics, fuel usage, tyre degradation and suspension geometry. This is because these have a negligible effect for road vehicles driven under normal conditions. Additionally, some of these simulators provide visually sparse environments that limit the application of computer vision techniques as part of an autonomous control stack.
@@
In contrast, Assetto Corsa (AC) is a customisable, racing-specific simulator that has a large collection of virtualised world-renowned race tracks and cars rendered to a high level of visual quality. AC’s inherent customisablity allows it to be used to simulate existing real-world autonomous racing leagues: Indy Autonomous Challenge [9], Autonomous Karting Series [10], Roborace [11] and A2RL [12]. Although the simulator itself is closed source, a mature modification community exists which have exposed internal methods of customisation and developed a range of additional assets such as vehicles, circuits, and visual improvements. AC is also regularly on sale for \$4USD [13] making it significantly more accessible compared to proprietary simulators used in motorsport. The desirability of AC has also been validated by its use in other work. In  [14] there is a section dedicated to validation of their control solution using AC and in the future, an interface would be published openly; this has yet to happen. Most recently [15], developed concurrently to AARK, details a similar interface for AC. Their interface was not released with the capability to: provide agents with visual sensor information or simulated INS readings, generate machine vision data, run on Unix systems and do not provide a modular full stack autonomous control solution.
@@
Enter the Adelaide Autonomous Racing Kit (AARK); an open source toolkit that enables research, development and validation of autonomous systems capable of operating vehicles pushed to their limits. AARK represents a feature rich, extensible manifestation of autonomous racing researchers’ desire to validate and compare autonomous systems in AC. Built around AC, AARK offers researchers a platform capable of generating data to train machine learning systems in addition to validating and comparing autonomous systems via a series of packages:
@@
  • AC Interface (ACI) - A harness that enables recording agent actions (human & autonomous), and enables autonomous systems to receive observations from and control vehicles in AC
@@
  • AC Reinforcement Learning (ACRL) - A gym environment for training reinforcement learning agents
The source code to each of these packages is available on Github https://github.com/orgs/Adelaide-Autonomous-Racing-Kit/repositories with accompanying documentation https://www.adelaideautonomous.racing/.
@@ III ADELAIDE AUTONOMOUS RACING KIT
The Adelaide Autonomous Racing Kit (AARK) positions itself in the no man’s land between two separate philosophical approaches to the autonomous racing problem; classical control and reinforcement learning (RL). Classical approaches to autonomous racing systems are built up from modular components that each handle specific tasks. This leads to systems that require laborious hand tuning of non-linear optimisation parameters which are specific to each particular vehicle and circuit [16, 17, 14, 18]. Conversely, RL approaches ingest environmental state and directly produce control outputs. In the current literature, classical autonomous systems heavily rely on known pose, GPS, LiDAR, and vehicle dynamics modelling. By contrast, previous RL methods access oracle knowledge of the simulation’s internal state [19, 20, 21, 15] and have the luxury of acting synchronously with their environment [22, 1]. For RL approaches to become practically useful, they need to evolve from relying on perfect state information and account for the fundamentally asynchronous nature of reality. The strengths of both approaches have the potential to address the weaknesses of the other. In fact, machine learning approaches are being integrated into classical autonomous system’s modules [23], albeit slowly, over time. For these fields to unify and develop together they require a common environment that is equally convenient for researchers of both methodologies to use. AARK has the potential to bridge this gap by providing a development and evaluation platform for autonomous racing systems that cater to both of these disciplines.
@@ IV ASSETTO CORSA INTERFACE
At its core AARK provides an interface between the popular racing simulator AC and autonomous agents, Figure 1. The enabling package, ACI, achieves this by scraping vehicle state from the simulator’s memory, capturing frames displayed using FFMPEG and relaying control inputs to AC via a virtual controller. At a frequency of 1̃00Hz, ACI outputs 190 different readings describing the simulated vehicle’s current state. These readings contain detailed information regarding the vehicle’s ground truth state, intended for evaluation and debugging. Accompanying this are 17 readings from a simulated Inertial Navigation System (INS), intended to be passed to the autonomous driver. The INS provides accelerometer, gyroscope, odometer and GPS readings with artificial noise and drift to better reflect real sensors. With these readings, ACI also provides image frames captured of the simulator’s application window. Frame captures are non-blocking, allowing systems to continue performing operations that depend on sensor readings while waiting for new visual input. The frequency of frame captures is 30Hz by default but can be configured by the user to suite their requirements and computer’s performance. Once an observation has been processed by the autonomous system it returns $\text{throttle},\text{brake},\text{steering}]$ to ACI which represents the inputs that should be applied to the vehicle.
Additionally, ACI provides two methods to capture driver data for analysis or training. Leveraging an AC plugin, AC Telemetry Interface (ACTI), vehicle performance data can be captured and written out in the prolific MoTeC format. MoTeC data is currently used to analyse both vehicle and driver performance across motorsport, providing access to a suite of ready-made visualisation tools to verify system behaviour. Vast libraries of driver MoTeC data exist online, for a wide range of operator proficiencies, making it an excellent way to compare autonomous and human driver behaviors. However, MoTeC does not offer a convenient way to package visual information synchronised with state information in the necessary form for training AI-based approaches. ACI addresses this deficit by recording sessions as a series of image, state pairs at a frequency of  30Hz. This format of recording can be used to train machine vision systems with, either directly or indirectly via post-processed data produced by AC Data Generator (ACDG). More details on the implementation framework for ACI are given in appendix IX-A.
@@ V DATA GENERATOR
Machine learning has demonstrated efficacy in tasks that estimate the properties of objects from images of those objects. Many of these properties can be leveraged to provide information to an autonomous system about its environment, such as depth, surface normal vectors and semantics. These sources of information have been under utilised in previous approaches that rely heavily on LiDAR, raw camera feeds, or ego state. The barrier, and likely reason for their lack of proliferation, to applying machine learning based vision techniques is their requirement for a large body of domain specific annotated data. For example, to train a semantic segmentation model for use in autonomous racing, video data from each different race track needs to be gathered and human annotators need to specify a label for each pixel in each image corresponding to a racing specific semantic class. ACDG automatically generates multiple types of annotated machine learning data from recordings gathered by ACI. Figure 2 shows an example of semantic, depth, and surface normal annotations generated for an image by ACDG. This is achieved by; recreating the recorded camera perspective at each time step, casting rays, and extracting information about where those rays intersect with the circuit’s mesh. Depth and surface normal annotations can be generated for any circuit.
ACDG’s initial release contains semantic mappings for eight circuits, shown in Figure 4. These include six official AC maps: Imola, Circuit de Spa-Francorchamps, Autodromo Nazionale di Monza, Silverstone, Vallenlunga, Nordschleife, in addition to two community-made tracks: Mount Panorama and Yas Marina. Tracks were chosen based on their respective challenge to drivers and agents, alike, with varying levels of complexity. The complexity of a track ranges from Monza, which has many long straights and very few corners, to Nurburgring - one of the most challenging tracks in the world, with 170 turns spaced over 25km. Mt Panorama and Spa feature large changes in elevation and blind corners. Silverstone and Monza contain several corner elements that need to be connected to prioritise exit speed onto straights rather than speed through corners. Vallelunga focuses on extremely technical slow speed cornering with hairpin style loops. Yas Marina and Imola are contain a variety of circuit elements and are included to support parity with real-world autonomous racing leagues, such as A2RL [12].
Each of these race tracks includes a semantic segmentation dataset and a pre-trained deep learning model for use in researchers’ control solutions. If researchers wish to develop a dataset for a not yet supported circuit, a series of video tutorials demonstrating how to add semantic segmentation data generation support for new tracks are available here: https://www.adelaideautonomous.racing/docs/acdg/getting-started/tutorials/.
@@ V-A DATASETS
To accelerate researcher’s ability to participate in autonomous racing, a series datasets created using AARK are available here: https://www.adelaideautonomous.racing/downloads/.
@@ V-A1 Behaviour
For each supported racetrack, associated recordings of human driving were gathered. Roughly eight laps of human control is available for use in both the ACI and MoTeC formats. MoTeC is compatible with i2, a free data analysis software used by professional motorsport teams to analyse driver and vehicle performance. By collecting telemetry data in this format, autonomously controlled vehicle behaviour can be compared against their human driver counter parts. This can provide insights into how autonomous control differs from race car drivers and vice versa. Recordings provided in the ACI format are compatible with behavioural cloning or supervised machine learning methods. Additionally, the ACI format can be used to generate a series of machine vision data for perception systems based on deep learning.